F2C for VAX/VMS

Date of last revision: June 4, 1990

This document describes the use of f2c, a filter for translating input files which are Fortran-77 code to either C or C++ code. This document may be printed on the line printer in CPE-3.102 by entering the command;

DOCPRINT SYS$DOCUMENT:F2C.DOC

Copies of this document may be printed on the laser printer in CPE-3.159 by entering the command;

PSPRINT SYS$DOCUMENT:F2C.PS

You may view copies of this document on workstations or personal computers which recognize the X window system protocol by entering the command;

XDVI SYS$DOCUMENT:F2C.DVI

Description

F2c converts Fortran 77 source code in file to C (or C++) a source file in the current directory, with .c substituted for the file type extension. If no Fortran files are named, f2c reads Fortran from standard input and writes C on standard output. File names that end with .P are taken to be prototype files, as produced by option -P, and are read first.

Format

f2c [option ...] file

f2c is compiled with VAXC which converts uppercase command line arguments to lowercase. To pass an uppercase argument or a negation argument to f2c, the argument must be enclosed in quotation marks ( i.e. “-C” or “-!bs”). Arguments enclosed in brackets ( [ ] ) are optional.

Options

The following options have the same meaning as in f77(1) on Unix systems.

-C Compile code to check that subscripts are within declared array bounds.

-I2 Render INTEGER and LOGICAL as short, INTEGER*4 as long int. Assume the default F2CLIB.OLB library: allow only INTEGER*4 (and no LOGICAL) variables in INQUIREs. Option -I4 confirms the default rendering of INTEGER as long int.

-onetrip Compile DO loops that are performed at least once if reached. (Fortran 77 DO loops are not performed at all if the upper limit is smaller than the lower limit.)

-U Honor the case of variable and external names. Fortran keywords must be in lower case.

-u Make the default type of a variable ‘undefined’ rather than using the default Fortran rules.

-w Suppress all warning messages. If the option is -w66, only Fortran 66 compatability warnings are suppressed.

The following options are peculiar to f2c. -A Produce ANSI C. Default is old-style C. -a Make local variables automatic rather than static unless they appear in a “DATA, EQUIVA-

LENCE, NAMELIST,” or SAVE statement. -C++ Output C++ code.

-c Include original Fortran source as comments. -E Declare uninitialized COMMON to be Extern (overridably defined in f2c.h as extern). -ec Place uninitialized COMMON blocks in separate files: COMMON /ABC/ appears in file abc com.c.

Option -e1c bundles the separate files into the output file, with comments that give an unbundling sed(1) script.

-ext Complain about f 77(1) extensions. -g Include original Fortran line numbers as comments. -h Try to align character strings on word (or, if the option is -hd , on double-word) boundaries. -i2 Similar to -I2, but assume a modified F2CLIB.OLB library (compiled with -Df2c i2 ), so INTE-

GER and LOGICAL variables may be assigned by INQUIRE. -kr Use temporary values to enforce Fortran expression evaluation where K&R (first edition) paren-

thesization rules allow rearrangement. If the option is -krd, use double precision temporaries even for single-precision operands.

-P Write a file.P of ANSI (or C++) prototypes for procedures defined in each input file.f or file.F. When reading Fortran from standard input, write prototypes at the beginning of standard output. Implies -A unless option -C++ is present. Option -Ps implies -P, and gives exit status 4 if rerunning f2c may change prototypes or declarations.

-p Supply preprocessor definitions to make common-block members look like local variables. -R Do not promote REAL functions and operations to DOUBLE PRECISION. Option -!R confirms

the default, which imitates f 77. -r8 Promote REAL to DOUBLE PRECISION, COMPLEX to DOUBLE COMPLEX.

-Tdir Put temporary files in directory dir. -w8 Suppress warnings when COMMON or EQUIVALENCE forces odd-word alignment of doubles. -Wn Assume n characters/word (default 4) when initializing numeric variables with character data.

-z Do not implicitly recognize DOUBLE COMPLEX. -!bs Do not recognize backslash escapes (\”, \’, \0, \\, \b, \f, \n, \r, \t, \v) in character strings. -!c Inhibit C output, but produce -P output. -!I Reject include statements. -!it Don’t infer types of untyped EXTERNAL procedures from use as parameters to previously defined

or prototyped procedures.

-!P Do not attempt to infer ANSI or C++ prototypes from usage.

The resulting C invokes the support routines of f 77; object code should be linked with the f2c support library as well as the VAXC run-time library. The appropriate LINK command under VMS would be;

$ LINK file,SYS$PUBLIC:F2CLIB/LIB,SYS$LIBRARY:VAXCRTL/LIB

Calling conventions are those of f77: see the reference below.

Files

file.[f][F][for][FOR] input file

*.c output file

SYS$PUBLIC:F2C.H header file

SYS$PUBLIC:F2CLIB.OLB intrinsic function library and Fortran I/O library

SYS$LIBRARY:VAXCRTL.OLB C run-time support library

See Also

S. I. Feldman and P. J. Weinberger, ‘A Portable Fortran 77 Compiler’, UNIX Time Sharing System Programmer’s Manual, Tenth Edition, Volume 2, AT&T Bell Laboratories, 1990.

You should also read the technical report on f2c from AT&T which provides a more detailed explanation of f2c. This file is normally found in SYS$DOC:F2C TECHNICAL REPORT.PS.

Diagnostics

The diagnostics produced by f2c are intended to be self-explanatory.

Bugs

Floating-point constant expressions are simplified in the floating-point arithmetic of the machine running f2c, so they are typically accurate to at most 16 or 17 decimal places.

Untypable EXTERNAL functions are declared int.